Consider the following two mutually exclusive projects:

  

Year Cash Flow (A) Cash Flow (B)
0 –$313,058        –$15,540         
1 29,800        5,197         
2 59,000        8,698         
3 58,000        13,394         
4 425,000        8,632         

  

Whichever project you choose, if any, you require a 6 percent return on your investment.
a. What is the payback period for Project A?

b. What is the payback period for Project B?

c. What is the discounted payback period for Project A?
Answer #1

a.  Payback Period Project A= ( Last Year with a Negative Cash Flow ) + [( Absolute Value of negative Cash Flow in that year)/ Total Cash Flow in the following year)]

= 3 + (166,258/ 425,000)

= 3.39 Years

Hence the correct answer is 3.39 Years

Note:

Year Investment Cash Inflow Net Cash Flow
0 -3,13,058 -    -3,13,058 (Investment + Cash Inflow)
1 -    29,800 -2,83,258 (Net Cash Flow + Cash Inflow)
2 -    59,000 -2,24,258 (Net Cash Flow + Cash Inflow)
3 -    58,000 -1,66,258 (Net Cash Flow + Cash Inflow)
4 -    4,25,000 2,58,742 (Net Cash Flow + Cash Inflow)

b. Payback Period Project B= ( Last Year with a Negative Cash Flow ) + [( Absolute Value of negative Cash Flow in that year)/ Total Cash Flow in the following year)]

= 2 + ( 1645 / 13394)

= 2.12 Years

Hence the correct answer is 2.12 Years

Note:

Year Investment Cash Inflow Net Cash Flow
0 -15,540 -    -15,540 (Investment + Cash Inflow)
1 -    5,197 -10,343 (Net Cash Flow + Cash Inflow)
2 -    8,698 -1,645 (Net Cash Flow + Cash Inflow)
3 -    13,394 11,749 (Net Cash Flow + Cash Inflow)
4 -    8,632 20,381 (Net Cash Flow + Cash Inflow)

c. Discounted Payback Period = ( Last Year with a Negative Cumulative Cash Flow ) + [( Absolute Value of negative Cumulative Cash Flow in that year)/ Total Present Cash Flow in the following year)]

= 3 + ( 183737.084076117 / 336639.806876159)

= 3.55 Years

Hence the correct answer is  3.55 Years

Note:

Cash Flow Discounting Factor ( 6%) Present Value (Cash Flow * Discounting Factor) Cumulative Cash Flow (Present Value of Current Year+ Cumulative Cash Flow of Previous Year)
0 -3,13,058 1 -3,13,058.00 -3,13,058.00
1 29,800 0.943396226415 28,113.207547170 -2,84,944.792453
2 59,000 0.889996440014 52,509.789960840 -2,32,435.002492
3 58,000 0.839619283032 48,697.918415874 -1,83,737.084076
4 4,25,000 0.792093663238 3,36,639.806876159 1,52,902.722800